The data below represents the number of T-shirts sold per week by a student who started his own online t-shirt business. Find the weighted mean of the number of t-shirts sold per week. (Round your answer to the nearest tenth if necessary.)
T-Shirts Sold per Week Frequency
4 1
8 7
12 5
16 6

Answer by mananth(16946)

4 1
8 7
12 5
16 6
(4* 1+ 8* 7+ 12* 5 +16* 6)=216
216/4 =54 ( 4 observations)